write down the value of :
cot310
tan(-180)
cos(-170)
sec(-170)
cosec(210)
tan(250)

Answer by Alan3354(69443)

write down the value of :
cot310
= 1/tan(310)
tan(-180)
= tan(180)
--------------
cos(-170)
= cos(170)
----------
sec(-170)
= 1/cos(170)
-----------
cosec(210)
= 1/sin(210)
-----------
tan(250)
Use a calculator.